OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (anonymous):

@Nnesha @Agl202

OpenStudy (agl202):

@ribhu

Nnesha (nnesha):

parallel lines have have same slope do you know what is slope of this equation\[\huge y = \frac{ 1 }{ 4 }x +2\]

Nnesha (nnesha):

remember y = mx + b this is slope intercept form equation where m is slope which is always with x and b is y-intercept so what is m in this equation y=1/4x+2

OpenStudy (ribhu):